OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

i assume he counted the horses heads as well as the humans so 74/2=37 humans 196/4= 49 horses, assuming they all have all of their legs

OpenStudy (anonymous):

that is wrong because we have to subtract the number of human legs from the horse's legs....hang on

OpenStudy (anonymous):

37 humans =74 legs to subtract. 196-74 =122 122/4=30.5 and that cannot be right is he including himself?